# Universal Modelling of Emergent Oscillations in Fractional Quantum Hall Fluids

### v1

@@ Abstract
To check the accuracy of the model, we compare the density profile of the quasihole in the thermodynamic limit produced by our method and the polynomial expansion method proposed in Ref. [29]. The two methods deviate very slightly from each other (root-mean-square difference is $2{\times}10^{-5}$) producing nearly coincident curves as shown in Fig. 2(a). This implies that, even though our model is very simple with only two free parameters, our result is as accurate as the polynomial expansion method which uses several tens of parameters.
$N_{e}$
$A_{1}$
$k_{1}$
$k_{1}*x_{1}$
$1/\lambda_{1}$
$R^{2}$
10
0.6757
1.4734
3.3342
0.8051
0.999883
11
0.6774
1.4646
3.3305
0.8097
0.999893
12
0.6787
1.4574
3.3277
0.8133
0.999904
13
0.6799
1.4514
3.3253
0.8164
0.999910
14
0.6809
1.4462
3.3231
0.8191
0.999915
15
0.6819
1.4417
3.3211
0.8214
0.999918
16
0.6827
1.4377
3.3194
0.8234
0.999921
17
0.6835
1.4343
3.3178
0.8252
0.999923
$\infty$
0.6944
1.3785
3.2947
0.8538
0.999982
Table S1: Parameters of the damped oscillator model for the density of a single Laughlin quasihole at $\nu{=}1/3$ for $N_{e}$ electrons in the spherical geometry. The statistical measure $R^{2}$ is the coefficient of determination for how well the modeled occupation predicts the exact occupation numbers. The final row is the thermodynamic extrapolation result as a function of $1/N_{e}$.
$N_{e}$
$A_{1}$
$k_{1}$
$k_{1}*x_{1}$
1/$\lambda_{1}$
$R^{2}$
10
2.1345
1.5208
4.8670
0.7670
0.999885
11
2.1446
1.5109
4.8520
0.7720
0.999886
12
2.1514
1.5028
4.8402
0.7759
0.999885
13
2.1571
1.4961
4.8304
0.7792
0.999887
14
2.1625
1.4902
4.8219
0.7821
0.999888
15
2.1676
1.4851
4.8143
0.7846
0.999890
16
2.1724
1.4806
4.8076
0.7869
0.999891
$\infty$
2.2335
1.4138
4.7094
0.8198
0.999900
Table S2: Similar to Table  S1 except the parameters are for $3{-}$Laughlin quasiholes at $\nu{=}1/3$.
$N_{e}$
$A_{1}$
$k_{1}$
$k_{1}*x_{1}$
1/$\lambda_{1}$
$R^{2}$
20
5.2491e-04
1.5436
12.5625
0.6400
0.999889
25
2.1314e-04
1.5339
14.4496
0.6442
0.999894
30
9.4669e-05
1.5271
16.1576
0.6474
0.999899
35
4.4788e-05
1.5220
17.7294
0.6503
0.999904
40
2.2664e-05
1.5181
19.1959
0.6516
0.999905
$\infty$
0
1.4929
$\infty$
0.6635
0.999922
Table S3: Parameters for the Laughlin edge with $N_{e}$ electrons at $\nu{=}1/3$. $R^{2}$ is the R-squared value for the real-space density obtained by the Monte Carlo method and the modeled real-space density in the planar disk geometry. In the thermodynamic limit, $A_{1}{=}0$ and $x_{1}{=}\infty$ because the edge is located at the position $r{=}\infty$. Values in the last row are the thermodynamic extrapolation results obtained from a linear fit of the parameters as a function of $1/N_{e}$.
$N_{e}$
$A_{1}$
$k_{1}$
$k_{1}*x_{1}$
1/$\lambda_{1}$
$R^{2}$
15
4.6540
1.1986
5.1067
0.4515
0.999970
20
4.7661
1.1839
5.0775
0.4595
0.999974
25
4.8301
1.1755
5.0613
0.4640
0.999976
30
4.8715
1.1701
5.0510
0.4669
0.999977
50
4.9478
1.1596
5.0317
0.4722
0.999979
$\infty$
5.0791
1.1425
4.9980
0.4815
0.999984
Table S4: Similar to Table  S3 except the parameters are for the ground state pair-correlation function at $\nu{=}1/5$.
SII 1/3 Laughlin state pair-correlation function
We repeat the single quasihole analysis for $3$-Laughlin quasiholes at $\nu{=}1/3$. The parameters are shown in Table  S2 and the finite size scaling is shown in Fig. S1. The ground state pair-correlation function $g(r)$ in the thermodynamic limit can be obtained by dividing the density $\rho(r)$ of $3$-Laughlin quasiholes by the average density $\bar{\rho}{=}1/(2\pi m)$. Then, one can use the modeled $g(r)$ to calculate the per-particle variational energy of the Laughlin state for various interactions $V(r)$ through $V{=}(\bar{\rho}/2)\int d^{2}\,\bm{r}V(r)[g(r){-}1]$. For the Coulomb interaction $V(r){=}1/r$, the calculated energy $V$ with our modeled $g(r)$ is ${-}0.4096$, which is very close to the extrapolated result ${-}0.4098$ obtained from the exact diagonalization calculation [34].
Figure S1: (a) Finite-size scaling of the wave vector $k_{1}$ for the single Laughlin quasihole (blue squares), $3$-Laughlin quasiholes (red circles), and the Laughlin edge (black hexagrams) at $\nu{=}1/3$. The lines are linear fits as a function of $1/N_{e}$. (b) Finite-size scaling of the inverse of the decay length $1/\lambda_{1}$.
SIII 1/3 Laughlin edge
We create a Laughlin edge by placing the Laughlin state on the disk geometry. Since the edge density fluctuation is large, we need a large system size (beyond the ones accessible to exact diagonalization) to reliably study it. Therefore, we determine the edge density using the Monte Carlo method. Then, we obtain the parameters of the damped model by fitting them to the real-space density. As shown in Fig. S2, the real-space density profile of the Laughlin edge can also be well-represented with the damped model in the occupation-number space. The parameters for different sizes are shown in Table  S3 and the finite size scaling is shown in Fig. S1.
Figure S2: The density distribution of $1/3$ Laughlin state for $N_{e}=40$ in the disk (blue line). The red crosses denote the fitting using our damped model.
SIV 1/5 Laughlin state pair-correlation function
We also study the ground state pair-correlation function for $\nu{=}1/5$ Laughlin state using the Monte Carlo method and determine the parameters of the damped model by fitting the model to the real-space profile. As shown in Fig. S3, the pair-correlation function for the 1/5 Laughlin state can also be well-represented by our damped model. The parameters for different sizes and their thermodynamic extrapolation are shown in Table  S4. Then, we use the modeled $g(r)$ to calculate the per-particle variational energy of the Laughlin state in terms of the Coulomb interaction $V(r){=}1/r$, and the result is ${-}0.3278$, which is very close to the extrapolated result ${-}0.3275(1)$ obtained from exact diagonalization [35].
Figure S3: The ground state pair-correlation function for the $1/5$ Laughlin state for $N_{e}{=}50$ electrons in the spherical geometry (blue line). The red crosses denote the results obtained using our damped model.
SV General cases
By using the Jack polynomial method or Monte Carlo method, we can determine the occupation-number density or real-space density of Laughlin quasihole for various $m$ and $n$. Then, we can determine their characteristic lengths as done above and the results are summarized in Fig. 3. Based on the numerical results for various Laughlin quasiholes, we find the damped oscillation model can always reasonably capture the density-distribution of quasiholes, but its numerical accuracy ($R^{2}$ value) decreases increasing $m$ and $n$. The numerical accuracy dependence on $n$ can be seen by comparing the $R^{2}$ value of Table  S1 for $n{=}1$ and Table  S2 for $n{=}3$. The numerical accuracy dependence on $m$ can be seen by comparing the fitting in Fig. 2 for $m{=}3$ and Fig. S4 for $m{=}10$. For large values of $m$ and $n$, there are stronger finite-size effects which can potentially explain the lower numerical accuracy of the damped model. For these large values of $m$ and $n$, it is also possible that our model is unable to capture the physical features of the states accurately. When the density fluctuation is very strong, especially for large $m$, it is better to determine the characteristic lengths of the oscillation tail by starting from a finite $r$, e.g., the first zero point of $\delta\rho(r){=}0$.
Figure S4: The real-space density of the single Laughlin quasihole at $\nu{=}1/10$ and $N_{e}{=}100$ electrons on the sphere (blue circles). The red crosses denote the results obtained using our damped model.
SVI Coulomb quasihole
In addition to the single-mode model of the Laughlin quasihole, another mode with four more fitting parameters is introduced for the Coulomb quasihole. The fitting parameters are shown in Table  S5. Note that numerical results show that there exist multiple sets of parameters that fit the density distribution very well. Therefore, it is possible that overfitting could be an issue for this model. We also find strong finite-size-scaling effects that hinder an accurate extrapolation to the thermodynamic limit.
$N_{e}$
$A_{1}$
$k_{1}$
$k_{1}*x_{1}$
$1/\lambda_{1}$
$A_{2}$
$k_{2}$
$k_{2}*x_{2}$
$1/\lambda_{2}$
$R^{2}$
10
0.2039
1.4965
2.7515
0.3778
0.5390
0.6411
2.0603
0.9143
0.999694
11
0.2283
1.4794
2.7356
0.4078
0.4948
0.6186
2.0792
0.8994
0.999920
12
0.2317
1.5039
2.8840
0.4202
0.5449
0.4572
1.5763
0.9318
0.999933
13
0.2232
1.4982
2.8624
0.4129
0.5375
0.5095
1.7350
0.9445
0.999845
14
0.2262
1.4847
2.8172
0.4160
0.5115
0.5425
1.8643
0.9318
0.999861
Table S5: Parameters of the double-damped model for the Coulomb quasihole with $N_{e}$ electrons at $\nu{=}1/3$.
$N_{e}$
$A_{1}$
$k_{1}$
$k_{1}*x_{1}$
$1/\lambda_{1}$
$A_{2}$
$k_{2}$
$k_{2}*x_{2}$
$1/\lambda_{2}$
$R^{2}$
15
0.9540
1.5628
5.7802
0.5936
7.3000
1.4715
3.4053
1.5951
0.999434
17
0.9761
1.5274
5.6332
0.6076
7.5191
1.4616
3.3479
1.6510
0.999540
19
1.0409
1.5138
5.5905
0.6273
6.9436
1.4817
3.3604
1.6286
0.999566
21
1.0593
1.5009
5.5481
0.6354
6.6887
1.4837
3.3545
1.6288
0.999598
$\infty$
1.3437
1.3457
4.9646
0.7451
5.0878
1.5220
3.2328
1.7069
Table S6: Parameters of the double-damped model for the Moore-Read quasihole density proportional to the ground state pair-correlation function. The last row is the thermodynamic extrapolation result as a function of $1/N_{e}$.
Figure S5: (a) The occupation numbers of the Moore-Read quasihole, the density of which is proportional to the Moore-Read ground state pair-correlation function, for $N_{e}{=}21$ electrons in the spherical geometry. Blue circles (red crosses) denote the exact (fitted) density. The inset is the corresponding real-space density. (b) The Moore-Read ground state pair-correlation function in the thermodynamic limit. The blue line and red crosses denote the results derived using the polynomial expansion method with $45$ basis [29] and our double-damped model, respectively. The inset is the zoomed view of the oscillation tail.
SVII pair-correlation function of Moore-Read state
The double-damped oscillation model also gives a good representation of the densities of Abelian quasiholes obtained from flux insertion (but no further fractionalization) for many other kinds of model FQH states. As an example, we show how to model the pair-correlation function of the Moore-Read ground state in the thermodynamic limit from it. To reduce the finite-size-scaling effects, we first determine the parameters of the long-range mode by fitting with the occupation numbers at a large $r$ (here our fitting starts from $n_{8}$), and then we determine the parameters of the short-range mode using all occupation numbers except the first two special orbitals. In this way, the finite-size effect is greatly reduced and meanwhile, the accuracy is still good as shown in Table  S6 and Fig. S5(a). The result obtained from an extrapolation to the thermodynamic limit is shown in Fig. S5(b). For comparison, we also plot the corresponding result obtained from the polynomial expansion method [29]. The two methods produce very similar results with a root-mean-square difference of $2{\times}10^{-3}$). Nevertheless, the density profile obtained from our model is smoother than that obtained using the polynomial expansion method.

### v2

@@ Appendix A FQH states and Jack polynomials
To check the accuracy of the model, we compare the density profile of the quasihole in the thermodynamic limit produced by our method and the polynomial expansion method proposed in Ref. [29]. The two methods deviate very slightly from each other (the root-mean-square difference is $2{\times}10^{-5}$) producing nearly coincident curves as shown in Fig. 2(a). This implies that, even though our model is very simple with only two free parameters, our result is as accurate as the polynomial expansion method which uses several tens of parameters.
$N_{e}$ $A_{1}$ $k_{1}$ $k_{1}*x_{1}$ $1/\lambda_{1}$ $R^{2}$
10 0.6757 1.4734 3.3342 0.8051 0.999883
11 0.6774 1.4646 3.3305 0.8097 0.999893
12 0.6787 1.4574 3.3277 0.8133 0.999904
13 0.6799 1.4514 3.3253 0.8164 0.999910
14 0.6809 1.4462 3.3231 0.8191 0.999915
15 0.6819 1.4417 3.3211 0.8214 0.999918
16 0.6827 1.4377 3.3194 0.8234 0.999921
17 0.6835 1.4343 3.3178 0.8252 0.999923
$\infty$ 0.6944 1.3785 3.2947 0.8538 0.999982
Table 1: Parameters of the damped oscillator model for the density of a single Laughlin quasihole at $\nu{=}1/3$ for $N_{e}$ electrons in the spherical geometry. The statistical measure $R^{2}$ is the coefficient of determination for how well the modeled occupation predicts the exact occupation numbers. The final row is the thermodynamic extrapolation result as a function of $1/N_{e}$.
Figure 7: (a) Finite-size scaling of the wave vector $k_{1}$ for the single Laughlin quasihole (blue squares), $3$-Laughlin quasiholes (red circles), and the Laughlin edge (black hexagrams) at $\nu{=}1/3$. The lines are linear fits as a function of $1/N_{e}$. (b) Finite-size scaling of the inverse of the decay length $1/\lambda_{1}$.
$N_{e}$ $A_{1}$ $k_{1}$ $k_{1}*x_{1}$ 1/$\lambda_{1}$ $R^{2}$
10 2.1345 1.5208 4.8670 0.7670 0.999885
11 2.1446 1.5109 4.8520 0.7720 0.999886
12 2.1514 1.5028 4.8402 0.7759 0.999885
13 2.1571 1.4961 4.8304 0.7792 0.999887
14 2.1625 1.4902 4.8219 0.7821 0.999888
15 2.1676 1.4851 4.8143 0.7846 0.999890
16 2.1724 1.4806 4.8076 0.7869 0.999891
$\infty$ 2.2335 1.4138 4.7094 0.8198 0.999900
Table 2: Similar to Table  1 except the parameters are for $3{-}$Laughlin quasiholes at $\nu{=}1/3$.
$N_{e}$ $A_{1}$ $k_{1}$ $k_{1}*x_{1}$ 1/$\lambda_{1}$ $R^{2}$
20 5.2491e-04 1.5436 12.5625 0.6400 0.999889
25 2.1314e-04 1.5339 14.4496 0.6442 0.999894
30 9.4669e-05 1.5271 16.1576 0.6474 0.999899
35 4.4788e-05 1.5220 17.7294 0.6503 0.999904
40 2.2664e-05 1.5181 19.1959 0.6516 0.999905
$\infty$ 0 1.4929 $\infty$ 0.6635 0.999922
Table 3: Parameters for the Laughlin edge with $N_{e}$ electrons at $\nu{=}1/3$. $R^{2}$ is the R-squared value for the real-space density obtained by the Monte Carlo method and the modeled real-space density in the planar disk geometry. In the thermodynamic limit, $A_{1}{=}0$ and $x_{1}{=}\infty$ because the edge is located at the position $r{=}\infty$. Values in the last row are the thermodynamic extrapolation results obtained from a linear fit of the parameters as a function of $1/N_{e}$.
$N_{e}$ $A_{1}$ $k_{1}$ $k_{1}*x_{1}$ 1/$\lambda_{1}$ $R^{2}$
15 4.6540 1.1986 5.1067 0.4515 0.999970
20 4.7661 1.1839 5.0775 0.4595 0.999974
25 4.8301 1.1755 5.0613 0.4640 0.999976
30 4.8715 1.1701 5.0510 0.4669 0.999977
50 4.9478 1.1596 5.0317 0.4722 0.999979
$\infty$ 5.0791 1.1425 4.9980 0.4815 0.999984
Table 4: Similar to Table  3 except the parameters are for the ground state pair-correlation function at $\nu{=}1/5$.
Appendix C 1/3 Laughlin state pair-correlation function
We repeat the single quasihole analysis for $3$-Laughlin quasiholes at $\nu{=}1/3$. The parameters are shown in Table  2 and the finite size scaling is shown in Fig. 7. The ground state pair-correlation function $g(r)$ in the thermodynamic limit can be obtained by dividing the density $\rho(r)$ of $3$-Laughlin quasiholes by the average density $\bar{\rho}{=}1/(2\pi m)$. Then, one can use the modeled $g(r)$ to calculate the per-particle variational energy of the Laughlin state for various interactions $V(r)$ through $V{=}(\bar{\rho}/2)\int d^{2}\,\bm{r}V(r)[g(r){-}1]$. For the Coulomb interaction $V(r){=}1/r$, the calculated energy $V$ with our modeled $g(r)$ is ${-}0.4096$, which is very close to the extrapolated result ${-}0.4098$ obtained from the exact diagonalization calculation [33].
Appendix D 1/3 Laughlin edge
We create a Laughlin edge by placing the Laughlin state on the disk geometry. Since the edge density fluctuation is large, we need a large system size (beyond the ones accessible to exact diagonalization) to reliably study it. Therefore, we determine the edge density using the Monte Carlo method. Then, we obtain the parameters of the damped model by fitting them to the real-space density. As shown in Fig. 8, the real-space density profile of the Laughlin edge can also be well-represented with the damped model in the occupation-number space. The parameters for different sizes are shown in Table  3 and the finite size scaling is shown in Fig. 7.
Figure 8: The density distribution of $1/3$ Laughlin state for $N_{e}=40$ in the disk (blue line). The red crosses denote the fitting using our damped model.
Appendix E 1/5 Laughlin state pair-correlation function
We also study the ground state pair-correlation function for $\nu{=}1/5$ Laughlin state using the Monte Carlo method and determine the parameters of the damped model by fitting the model to the real-space profile. As shown in Fig. 9, the pair-correlation function for the 1/5 Laughlin state can also be well-represented by our damped model. The parameters for different sizes and their thermodynamic extrapolation are shown in Table  4. Then, we use the modeled $g(r)$ to calculate the per-particle variational energy of the Laughlin state in terms of the Coulomb interaction $V(r){=}1/r$, and the result is ${-}0.3278$, which is very close to the extrapolated result ${-}0.3275(1)$ obtained from exact diagonalization [34].
Figure 9: The ground state pair-correlation function for the $1/5$ Laughlin state for $N_{e}{=}50$ electrons in the spherical geometry (blue line). The red crosses denote the results obtained using our damped model.
$N_{e}$ $A_{1}$ $k_{1}$ $k_{1}*x_{1}$ $1/\lambda_{1}$ $A_{2}$ $k_{2}$ $k_{2}*x_{2}$ $1/\lambda_{2}$ $R^{2}$
10 0.2039 1.4965 2.7515 0.3778 0.5390 0.6411 2.0603 0.9143 0.999694
11 0.2283 1.4794 2.7356 0.4078 0.4948 0.6186 2.0792 0.8994 0.999920
12 0.2317 1.5039 2.8840 0.4202 0.5449 0.4572 1.5763 0.9318 0.999933
13 0.2232 1.4982 2.8624 0.4129 0.5375 0.5095 1.7350 0.9445 0.999845
14 0.2262 1.4847 2.8172 0.4160 0.5115 0.5425 1.8643 0.9318 0.999861
Table 5: Parameters of the double-damped model for the Coulomb quasihole with $N_{e}$ electrons at $\nu{=}1/3$.
$N_{e}$ $A_{1}$ $k_{1}$ $k_{1}*x_{1}$ $1/\lambda_{1}$ $A_{2}$ $k_{2}$ $k_{2}*x_{2}$ $1/\lambda_{2}$ $R^{2}$
15 0.9540 1.5628 5.7802 0.5936 7.3000 1.4715 3.4053 1.5951 0.999434
17 0.9761 1.5274 5.6332 0.6076 7.5191 1.4616 3.3479 1.6510 0.999540
19 1.0409 1.5138 5.5905 0.6273 6.9436 1.4817 3.3604 1.6286 0.999566
21 1.0593 1.5009 5.5481 0.6354 6.6887 1.4837 3.3545 1.6288 0.999598
$\infty$ 1.3437 1.3457 4.9646 0.7451 5.0878 1.5220 3.2328 1.7069
Table 6: Parameters of the double-damped model for the Moore-Read hole density proportional to the ground state pair-correlation function. The last row is the thermodynamic extrapolation result as a function of $1/N_{e}$.
Appendix F General Laughlin cases
By using the Jack polynomial method or Monte Carlo method, we can determine the occupation-number density or real-space density of Laughlin quasihole for various $m$ and $n$. Then, we can determine their characteristic lengths as done above and the results are summarized in Fig. 3. Based on the numerical results for various Laughlin quasiholes, we find the damped oscillation model can always reasonably capture the density-distribution of quasiholes, but its numerical accuracy ($R^{2}$ value) decreases increasing $m$ and $n$ as shown in Fig. 10. For large values of $m$ and $n$, there are stronger finite-size effects which can potentially explain the lower numerical accuracy of the damped model. For these large values of $m$ and $n$, it is also possible that our model is unable to capture the physical features of the states accurately. When the density fluctuation is very strong, especially for large $m$, it is better to determine the characteristic lengths of the oscillation tail by starting from a finite $r$, e.g., the first zero point of $\delta\rho(r){=}0$.
Figure 10: (a) $R^{2}$ value for Laughlin $n$-quasihole state at $\nu=1/3$ and $N_{e}=15$. The statistical measure $R^{2}$ is the coefficient of determination for how well the modeled occupation predicts the exact occupation numbers obtained by Jack polynomial. (b) $R^{2}$ value for Laughlin single quasiholes at $\nu=1/m$ and $N_{e}=12$.
Appendix G Coulomb quasihole
In addition to the single-mode model of the Laughlin quasihole, another mode with four more fitting parameters is introduced for the Coulomb quasihole. The fitting parameters are shown in Table  5. Note that numerical results show that there exist multiple sets of parameters that fit the density distribution very well. Therefore, it is possible that overfitting could be an issue for this model. We also find strong finite-size-scaling effects that hinder an accurate extrapolation to the thermodynamic limit.
For the correction to the Coulomb potential, we consider the Yukawa interaction $V(r){=}e^{{-}\lambda r}/r$. The overlap of the quasihole state with finite decay length $\lambda{=}l_{B}$ and the bare Coulomb quasihole is $0.9910$. Their density profiles are very close to each other. The double-damped model also works in this case as shown in Fig. 11. If we further increase $\lambda$, the quasihole state will approach the Laughlin quasihole gradually, in which case our model also works. Thus, we expect the double-damped model also works in the case of Yukawa interaction in the LLL.
Figure 11: The occupation numbers of the quasihole for $N_{e}{=}13$ electrons and $N_{\phi}{=}37$ fluxes in the spherical geometry. Blue circles (red crosses) denote the exact (fitted) density generated by ED with Yukawa interaction $V(r){=}e^{{-}r}/r$. The inset is the corresponding real-space density.
Appendix H pair-correlation function of Moore-Read state
As an example, we show how to model the pair-correlation function of the Moore-Read ground state in the thermodynamic limit from our double-damped model. To reduce the finite-size-scaling effects, we first determine the parameters of the long-range mode by fitting with the occupation numbers at a large $r$ (here our fitting starts from $n_{8}$), and then we determine the parameters of the short-range mode using all occupation numbers except the first two special orbitals. In this way, the finite-size effect is greatly reduced and meanwhile, the accuracy is still good as shown in Table  6 and Fig. 12(a). The result obtained from an extrapolation to the thermodynamic limit is shown in Fig. 12(b). For comparison, we also plot the corresponding result obtained from the polynomial expansion method [29]. The two methods produce very similar results with a root-mean-square difference of $2{\times}10^{-3}$). Nevertheless, the density profile obtained from our model is smoother than that obtained using the polynomial expansion method.
Figure 12: (a) The occupation numbers of the Moore-Read hole state, the density of which is proportional to the Moore-Read ground state pair-correlation function, for $N_{e}{=}21$ electrons and $N_{\phi}{=}41$ fluxes in the spherical geometry. Blue circles (red crosses) denote the exact (fitted) density generated by Jack polynomials with the root $010011...0011$. The inset is the corresponding real-space density. (b) The Moore-Read ground state pair-correlation function in the thermodynamic limit. The blue line and red crosses denote the results derived using the polynomial expansion method with $45$ basis [29] and our double-damped model, respectively. The inset is the zoomed view of the oscillation tail.